# Descriptive Image Quality Assessment in the Wild

### v2

@@ 3.3 Dataset Construction
High-quality and large-scale datasets are crucial for training VLMs. Following [26, 70], training VLMs requires {images, question, response} triplets, where “images” are the ones to be evaluated, “question” describes the task, and “response” is the ground truth answer. In this section, we detail the construction of our dataset from the selection of images and the collection of questions and responses.
@@
Image collection. Typical IQA datasets involve two types of images: high-quality reference images and distorted images to be evaluated. Generating distorted images is easy given our comprehensive distortion library introduced in Sec. 3.2. Existing studies often collect a large number of distorted images from a small number of references [15, 43]. However, the semantic richness of images is also crucial for VLM training. Therefore, we primarily source reference images from the KADIS-700K dataset [25], which offers 140K pristine reference images from diverse natural and daily scenes. We also leverage other IQA datasets for their convenience to generate responses (details in the following).
@@
Question collection. Humans often express similar questions using different sentences, necessitating model robustness to various user questions. For each task, we initially prompt GPT-4 [40] to generate 50 candidate questions. Subsequently, we manually eliminate ambiguous and repetitive ones and correct inaccurate ones, creating a question set of 20 questions (see Appendix). These questions are randomly sampled during training and testing to form the data pair.
@@
Response collection. We employ two response types as shown in Fig. 3. The first comprises brief templated responses that are easy to produce, where we emphasize the quantity to bring robust fundamental skills. The second consists of detailed responses, where we emphasize the quality to enhance the model’s advanced reasoning abilities. Existing methods to collect detailed responses mainly rely on human annotation [62, 71] and GPT-4V generation [64]. However, human annotation can be biased and vary in quality particularly when annotators are untrained or tired [71]. Also, GPT-4V is not fully reliable since its IQA performance is still unsatisfactory as evidenced in Sec. 5.
@@
We rethink the key aspects of our desired responses and GPT-4V’s corresponding abilities, introducing GT-informed generation by prompting the Ground Truth (GT) details to enhance GPT-4V’s generation. Specifically, a high-quality detailed response should contain image contents, key distortions, the impacts of distortions on contents, and conclusions (e.g., comparison results). While GPT-4V excels at identifying contents and analyzing impacts, it struggles with distortion identification and quality comparison, which will be shown in Sec. 5. To compensate for that, we directly provide it with explicit GT information. The response generation for each task is detailed subsequently.
@@
Figure 3: Construction of our DQ-495K dataset. For distortion identification, templated responses are generated using distortion information. In instant rating, we sample images from existing datasets and compare the Mean Opinion Score (MOS) to determine the better image for templated response creation. For assessment reasoning and comparison reasoning tasks, we provide GPT-4V with evaluated images and Ground Truth (GT) details (i.e., distortion information, comparison results from an assistant model) to facilitate detailed and accurate response generation, called GT-informed generation. This additional information is critical as GPT-4V cannot produce it accurately.
@@
Task1-brief: distortion identification. As shown in Fig. 3a, we first establish a response pool containing 20 templates with unspecified distortions. Next, we add distortions into the reference to create its distorted counterpart and populate a sampled template with the specific distortions to complete the response. For streamlined evaluation, inspired by [27], we randomly select half of the questions and append the short answer prompt: “Answer the question using a single word or phrase.” Correspondingly, the response will be a single phrase, like “noise”, specifying the distortions.
@@
Task1-detail: assessment reasoning. Given the reference image, we initially introduce distortions to corrupt the reference. Then, GPT-4V is input with both two images and the distortion information, and requested to assess the quality of the distorted image, as illustrated in Fig. 3b. We instruct GPT-4V to respond from three dimensions: contents, distortions along with their impacts on contents, and overall quality. Here prior studies [62, 64] primarily focus on low-level properties, while we consider how these low-level distortions influence the display of high-level contents.
@@
Task2-brief: instant rating. We begin by sampling a reference image and its two distorted versions from existing IQA datasets, and then compare the Mean Opinion Score (MOS) to determine the better one, as shown in Fig. 3c. Similar to distortion identification, we assemble a response pool of 20 templates to convert the comparison results into textural responses, and append the short answer prompt for the convenience of evaluation. We select three IQA datasets for training, including BAPPS [77], KADID-10K [24], and PIPAL [15], to cover a diverse range of reference images.
@@
Task2-detail: comparison reasoning. As depicted in Fig. 3d, given a high-quality image, we randomly apply distortions to produce two distorted images. We first train an assistant model using the large-scale instant rating data to predict the comparison results. Note that GPT-4V does not perform well on the quality comparison task, as shown in our experiments in Tab. 4 and Tab. 7, and thus we train our own comparison model. Then, similar to assessment reasoning, we inform GPT-4V of the three images, distortion information, and comparison results to generate detailed responses.
@@
Table 2: Statistics of our DQ-495K dataset.
Task1-brief Task1-detail Task2-brief Task2-detail
@@
Distortion Identification Assessment Reasoning Instant Rating Comparison Reasoning
Train 112,000 / 112,000 19,829 / 7,981 215,676 19,809 / 7,958
Validation 28,000 / 28,000 200 / 100 41,120 200 / 100
@@
Dataset statistics. The dataset statistics are illustrated in Tab. 2 (more in Appendix). All tasks, except instant rating, are displayed in the single-distortion / multi-distortion format. Our training set contains 439,676 brief samples and 55,577 detailed samples. For instant rating, the training set includes BAPPS, KADID, and PIPAL, while the validation set consists of BAPPS, KADID, PIPAL, TID2013 [43], LIVE-MD [17], and MDID2013 [53]. To ensure no intersection between training and validation sets for those overlapped datasets, the original splits are kept. For detailed tasks, all samples in the validation set have been carefully checked by humans.
4 Model Design
We primarily follow DepictQA [71] to construct and train our model. The base large language model is Vicuna-v1.5-7B [8]. Following [71], we also adopt the unique tag technique, i.e., using specific tags for various types of images, to ensure that the language model can distinguish different input images. To increase robustness, an external high-level perception dataset (detailed description dataset in [70]) is incorporated during training as a regularization, similar to [62, 71]. See details in Appendix.
@@ 4 Model Design
Retaining resolution in training. Although previous VLM-based IQA models typically resize all input images to a fixed resolution [62, 64], we find this might hurt their performance, as resolution variation may affect visual quality. Instead, we retain the original image resolution during training. Specifically, we interpolate (in bicubic mode) the position embedding in CLIP [45] to accommodate varying image resolutions. Ablation studies detailed in Sec. 5.4 demonstrate our model’s capability to assess quality variations attributable to resolution, even without explicitly training on such tasks.
@@
Confidence estimation. In many applications, it is important to know a confidence score that indicates when the model is uncertain of its response. Here we use the confidence scores of some key tokens as the confidence of the entire answer. Intuitively, the key tokens are distortion names in distortion identification, and are either “Image A” or “Image B” in instant rating. For detailed reasoning tasks, which feature diverse and non-structured responses, we utilize semantic change testing [12] to identify the top 20 tokens with the highest importance scores as key tokens. In semantic change testing, we employ all-MiniLM-L6-v2 [46] as the similarity model, due to its high processing speed (14K sentences per second). The predicted likelihood of key tokens is averaged as the confidence score. Fig. 6 verifies that confidence and model performance are highly correlated.
5 Experiments
We conduct extensive experiments on benchmarks and real-world applications including assessing web-downloaded images and comparing model-processed images, validating our advantages.
5.1 Metrics and Baselines
@@
Accuracy. The accuracy metric is utilized for distortion identification and instant rating tasks. VLMs usually produce diverse textual outputs, and we transform them into brief results for accuracy calculation. Specifically, we prompt our DepictQA-Wild with “Answer the question using a single word or phrase” to encourage direct output of brief responses. For baseline models, we include all potential answers in the prompt and instruct the model to identify the most accurate one.
@@
GPT-4 score as evaluation metric. We employ the GPT-4 score to evaluate assessment reasoning and comparison reasoning tasks, following [26]. Specifically, we provide GPT-4 with both the model-generated response and the corresponding ground truth response. GPT-4 assesses the helpfulness, relevance, accuracy, and level of detail in the model-generated response relative to the ground truth, assigning an overall score on a scale of 0 to 10, where a higher score indicates better quality. This average score is subsequently normalized to a scale of 0 to 100%, reported as the GPT-4 score metric.
@@
Baselines. We categorize baseline methods into general-purpose VLMs and IQA-specific VLMs. For general VLMs, we include mPLUG-Owl2 [68] (based on LLaMA-2-7B [56]), LLaVA-1.6 [28] (based on Vicuna-v1.5-7B [8]), and the proprietary GPT-4V [41]. IQA-specific VLMs are represented by Q-Instruct [62] and Co-Instruct [64]. Additionally, we compare traditional score-based IQA methods including full-reference ones (PSNR, SSIM [59], LPIPS [77], DISTS [10]) and non-reference ones (NIQE [37], ClipIQA [58], MUSIQ [20], MANIQA [66]) in instant rating task.
5.2 Results on Benchmarks
@@
Table 3: Distortion identification results under both single-distortion and multi-distortion cases. The accuracy metric is reported in the full-reference / non-reference settings. DepictQA-Wild greatly outperforms all baselines and maintains its high accuracy in out-of-distribution (OOD) setting.
@@
mPLUG-Owl2 [68] LLaVA-1.6 [28] GPT-4V [41] Q-Instruct [62] Co-Instruct [64] DepictQA-Wild DepictQA-Wild (OOD)
Single-dist. 10.1 / 11.6 14.0 / 15.3 46.7 / 45.2 12.9 / 15.5 27.2 / 34.4 97.9 / 94.7 82.1 / 73.2
Multi-dist. 10.8 / 10.7 12.0 / 12.1 41.5 / 39.8 21.5 / 23.9 30.2 / 33.3 90.5 / 89.5 76.6 / 77.2
Quantitative results of distortion identification are shown in Tab. 3. First, the performance of Q-Instruct and Co-Instruct is stably superior in the non-reference setting compared to the full-reference setting, attributed to their training without reference. Second, the performance of open-source general VLMs, including mPLUG-Owl2 and LLaVA-1.6, is still limited, but the proprietary GPT-4V [41] outperforms other general-purpose VLMs and exceeds prior specialized IQA VLMs. Third, DepictQA-Wild significantly surpasses all baseline methods, demonstrating our model’s efficacy. Finally, we evaluate our model in an out-of-distribution (OOD) setting. Specifically, for a particular category of distortion (e.g., noise), we use some sub-categories (e.g., Gaussian noise) during training, and different sub-categories (e.g., speckle noise) for evaluation. Results in the last column of Tab. 3 show that our method maintains high accuracy even under such an OOD setting.
Quantitative results of instant rating are demonstrated in Tab. 4. First, in the full-reference context, traditional score-based methods, even the simplest PSNR, outperform all general VLMs including GPT-4V and prior IQA-specific VLMs, indicating the inadequacy of existing VLMs in full-reference IQA tasks. Second, conversely, in the non-reference scenario, GPT-4V and Co-Instruct excel beyond most score-based approaches, except MUSIQ. Third, Co-Instruct is trained on multi-image comparison tasks without reference, and thus its performance in full-reference setting drops by quite a large margin. This further demonstrates the necessity of unifying full-reference and non-reference settings. Finally, DepictQA-Wild demonstrates superior performance across both settings by a large margin, showcasing its substantial advantage.
@@
Table 4: Instant rating results on multiple benchmarks in the full-reference / non-reference setting with the accuracy metric. DepictQA-Wild surpasses all baselines by a large margin.
Methods BAPPS [77] KADID [24] PIPAL [15] TID2013 [43] LIVE-MD [17] MDID2013 [53] Mean
Full-refer. Score-based IQA PSNR 68.9 / 78.7 / 80.9 / 85.0 / 89.7 / 78.0 / 80.2 /
SSIM [59] 69.7 / 77.1 / 82.6 / 78.7 / 88.1 / 76.8 / 78.8 /
LPIPS [77] 79.4 / 79.7 / 84.2 / 86.6 / 91.3 / 85.4 / 84.4 /
DISTS [10] 79.7 / 85.8 / 84.6 / 87.0 / 93.1 / 88.5 / 86.5 /
Non-refer. Score-based IQA NIQE [37] / 49.9 / 66.9 / 59.7 / 65.0 / 86.9 / 82.2 / 68.4
ClipIQA [58] / 59.7 / 75.8 / 72.6 / 85.8 / 65.8 / 47.0 / 67.8
MUSIQ [20] / 59.2 / 76.1 / 77.8 / 80.1 / 87.2 / 81.1 / 76.9
MANIQA [66] / 54.9 / 68.4 / 79.2 / 77.3 / 75.4 / 63.5 / 69.8
General VLM mPLUG-Owl2 [68] 50.1 / 50.1 50.6 / 50.8 49.6 / 49.6 48.6 / 48.5 49.9 / 50.1 50.6 / 50.5 49.9 / 49.9
LLaVA-1.6 [28] 54.1 / 56.2 50.4 / 51.9 52.0 / 52.6 54.2 / 57.0 54.4 / 56.5 54.3 / 53.1 53.2 / 54.6
GPT-4V [41] 70.3 / 63.2 83.2 / 81.5 78.5 / 78.2 84.4 / 88.1 79.6 / 72.7 70.6 / 67.6 77.8 / 75.2
IQA-specific VLM Q-Instruct [62] 50.4 / 50.1 51.8 / 55.8 50.4 / 51.9 50.8 / 55.0 51.8 / 53.0 49.6 / 49.6 50.8 / 52.6
Co-Instruct [64] 49.8 / 50.7 52.0 / 82.4 50.6 / 72.5 59.3 / 85.0 50.0 / 70.3 50.0 / 58.0 52.0 / 69.8
DepictQA-Wild 82.7 / 81.4 92.7 / 92.4 89.2 / 88.8 96.2 / 95.9 92.1 / 91.9 89.1 / 88.4 90.3 / 89.8

### v3

@@ III-C Dataset Construction
High-quality and large-scale datasets are crucial for training VLMs. Following [llava, lamm], training VLMs requires {images, question, response} triplets, where “images” are the ones to be evaluated, “question” describes the task, and “response” is the ground truth answer. In this section, we detail the construction of our dataset from the selection of images and the collection of questions and responses.
@@
Image collection. Typical IQA datasets involve two types of images: high-quality reference images and distorted images to be evaluated. Generating distorted images is easy given our comprehensive distortion library introduced in Sec. III-B. Existing studies often collect a large number of distorted images from a small number of references [pipal, tid2013]. However, the semantic richness of images is also crucial for VLM training. Therefore, we primarily source reference images from the KADIS-700K dataset [kadis], which offers 140K pristine reference images from diverse natural and daily scenes. We also leverage other IQA datasets for their convenience to generate responses (details are below).
@@
Question collection. Humans often express similar questions using different sentences, necessitating model robustness to various questions. For each task, we initially prompt GPT-4 [gpt4] to generate 50 candidate questions. Then, we manually eliminate ambiguous and repetitive ones and correct inaccurate ones, creating a set of 20 questions. We randomly sample these questions during training and testing to form the data pair.
@@
Response collection. We employ two response types as shown in Fig. 3. The first comprises brief templated responses that are easy to produce, where we emphasize the quantity to bring robust fundamental skills. The second consists of detailed responses, where we emphasize the quality to enhance the advanced reasoning abilities. Existing methods to collect detailed responses mainly rely on human annotation [qinstruct, depictqa] and GPT-4V generation [coinstruct]. However, human annotation can be biased and vary in quality when annotators are untrained or tired [depictqa]. Also, GPT-4V is not fully reliable since its IQA performance is still unsatisfactory as evidenced in Sec. V.
@@
We rethink the key aspects of our desired responses and GPT-4V’s corresponding abilities, introducing GT-informed generation by prompting the Ground Truth (GT) details to enhance GPT-4V’s generation. Specifically, a high-quality detailed response should contain image contents, key distortions, the impacts of distortions on contents, and conclusions. While GPT-4V excels at identifying contents and analyzing impacts, it struggles with distortion identification and quality comparison, as shown in Sec. V. To compensate for that, we directly provide it with explicit GT information. The response generation for each task is detailed subsequently.
@@
Figure 3: Construction of DQ-495K dataset. For distortion identification, templated responses are generated using distortion information. In instant rating, we sample images from existing datasets and compare the Mean Opinion Score (MOS) to determine the better image for templated response creation. For assessment reasoning and comparison reasoning tasks, we provide GPT-4V with evaluated images and Ground Truth (GT) details (i.e., distortion information, comparison results from an assistant model) to facilitate detailed and accurate response generation, called GT-informed generation. This additional information is critical as GPT-4V cannot produce it accurately.
@@
Task1-brief: distortion identification. As shown in Fig. 3a, we first establish a response pool containing 20 templates with unspecified distortions. Next, we add distortions into the reference to create its distorted counterpart and populate a sampled template with the specific distortions as the response. For streamlined evaluation, we sample half of the questions and append the short answer prompt: “Answer the question using a single word or phrase.” Correspondingly, the response will be a single phrase, like “noise”, specifying the distortions.
@@
Task1-detail: assessment reasoning. Given the reference image, we initially introduce distortions to corrupt the reference. Then, GPT-4V is input with both two images and the distortion information, and requested to assess the quality of the distorted image, as illustrated in Fig. 3b. We instruct GPT-4V to respond from three dimensions: contents, distortions along with their impacts on contents, and overall quality. Prior works [qinstruct, coinstruct] mainly focus on low-level properties, while we consider how these distortions influence the display of high-level contents.
@@
Task2-brief: instant rating. We begin by sampling a reference image and its two distorted versions from existing IQA datasets, and then compare the Mean Opinion Score (MOS) to determine the better one, as shown in Fig. 3c. Similar to distortion identification, we assemble a response pool of 20 templates to convert the comparison results into textual responses, and append the short answer prompt for the convenience of evaluation. We select three IQA datasets for training, including BAPPS [lpips], KADID-10K [kadid], and PIPAL [pipal], to cover a diverse range of reference images.
@@
Task2-detail: comparison reasoning. As depicted in Fig. 3d, given a high-quality image, we randomly apply distortions to produce two distorted images. We first train an assistant model using the large-scale instant rating data to predict the comparison results. Note that GPT-4V does not perform well on the quality comparison task, as shown in our experiments in Sec. V-B and Tab. IX, and thus we train our own comparison model. Then, similar to assessment reasoning, we inform GPT-4V of the three images, distortion information, and comparison results to generate detailed responses.
@@
TABLE II: Statistics of our DQ-495K dataset.
(a) Number of samples. All tasks except instant rating are displayed in the single-distortion / multi-distortion format. Task1-brief Task1-detail Task2-brief Task2-detail Distortion Identification Assessment Reasoning Instant Rating Comparison Reasoning Train 112,000 / 112,000 19,829 / 7,981 215,676 19,809 / 7,958 Test 28,000 / 28,000 200 / 100 41,120 200 / 100
@@
(b) Response length reported as word count / string length. Distortion Identification Assessment Reasoning Instant Rating Comparison Reasoning Single-dist. 10.36 / 69.81 64.37 / 430.23 9.30 / 52.02 93.20 / 604.97 Multi-dist. 12.84 / 88.67 87.31 / 588.44 114.04 / 740.68
III-D Dataset Analysis
@@
Dataset statistics are given in Tab. IIa. Our training set contains 439,676 brief samples and 55,577 detailed samples. For instant rating, the training set includes BAPPS, KADID, and PIPAL, while the validation set consists of BAPPS, KADID, PIPAL, TID2013 [tid2013], LIVE-MD [livemd], and MDID2013 [mdid]. To ensure there is no intersection between training and validation sets for those overlapped datasets, the original splits are kept. For detailed tasks, all samples in the validation set have been carefully checked by humans.
Figure 4: Statistics of all images in our dataset about (a) semantic features, (b) brightness, (c) contrast, (d) colorfulness, (e) edge density, and (f) texture variance.
Figure 5: Word length distribution of detailed responses.
Figure 6: Wordcloud map of all detail responses.
Image analysis. We analyze six statistical features to examine the image distributions. First, we use CLIP [clip] ViT to extract semantic features and apply K-Means clustering with $k=20$. As shown in Fig. 4a, the samples are relatively balanced across clusters. Second, the brightness (Fig. 4b) and contrast (Fig. 4c) approximately follow Gaussian distributions, consistent with natural image statistics [bright_contrast]. Third, as in Fig. 4de, colorfulness and edge density exhibit right-skewed heavy-tailed distributions, which are characteristics of natural scenes [color, edge_density]. Finally, texture variance (Fig. 4f) follows a highly right-skewed $\gamma$ distribution, aligning with prior findings [texture_variance]. These results confirm that our dataset maintains natural image statistics in semantic and low-level dimensions.
Text analysis. First, we examine the response length statistics of our DQ-495K dataset in Tab. IIb, including both word count and character length. For the instant rating task, single- and multi-distortion cases are not distinguished. We further present the word-length distribution of detailed reasoning responses in Fig. 6. The results show a near-Gaussian distribution centered around approximately 90 words, indicating that most responses are concise yet sufficiently descriptive. Second, Fig. 6 illustrates the word cloud of our DQ-495K dataset. We manually remove “Image A” and “Image B”, as they are constant proper nouns across all texts. The most frequent terms (e.g., “overly high”, “color quantization”, “high contrast”, “high saturation”, and “detail”) are closely related to low-level image attributes and visual quality.
IV Model Design
Model Architecture. DepictQA-Wild primarily adopts the architecture from LLaVA-1.6 [llavanext] and mPLUG-Owl2 [mplug-owl2], structured as follows. Specifically, the input images and the question texts are first tokenized, then fused, finally processed by the Large Language Model (LLM) for response generation. (1) Tokenizing input images and question texts. We use a frozen CLIP pre-trained ViT-L/14 [clip] as the image encoder to convert the input images into visual tokens. The question texts are tokenized into textual tokens using the SentencePiece tokenizer [sentencepiece]. To bridge the different embedding spaces of visual and textual tokens, we implement a trainable image abstractor, which is a four-layer transformer network, to map visual tokens into the textual space following [mplug-owl2]. The abstractor can also significantly reduce the number of vision tokens, relieving the computing pressure. (2) Token fusion. We integrate the visual tokens into pre-defined positions within the textual tokens as token fusion. (3) Response generation using LLM. The fused tokens are fed into LLM to generate the final response. Here we mainly conduct experiments with Vicuna-v1.5-7B. Despite their capabilities, pre-trained LLMs typically do not perform well on IQA tasks without adjustments. Therefore, we employ LoRA [lora], a fine-tuning technique that efficiently modifies a small subset of parameters within the LLM. Specifically, we apply LoRA to adjust the projection layers in all self-attention modules, following [lora, lamm]. This approach allows for targeted refinement of the model’s performance on IQA tasks without the need for extensive retraining. Our experiments in Tab. XIII also show that model architecture has little influence on model performance.
@@ IV Model Design
Retaining resolution in training. Although previous VLM-based IQA models typically resize all input images to a fixed resolution [qinstruct, coinstruct], we find this might hurt their performance, as resolution variation may affect visual quality. Instead, we retain the original image resolution during training. Specifically, we interpolate (in bicubic mode) the position embedding in CLIP [clip] to accommodate varying image resolutions. Ablation studies detailed in Sec. V-D demonstrate our model’s capability to assess quality variations attributable to resolution, even without explicitly training on such tasks.
@@
Confidence estimation. In many applications, it is important to know a confidence score that indicates when the model is uncertain of its response. Here we use the confidence scores of some key tokens as the confidence of the entire answer. Intuitively, the key tokens are distortion names in distortion identification, and are either “Image A” or “Image B” in instant rating. For detailed reasoning tasks, which feature diverse and non-structured responses, we utilize semantic change testing [sar] to identify the top 20 tokens with the highest importance scores as key tokens. In semantic change testing, we employ all-MiniLM-L6-v2 [sentence_transformer] as the similarity model, due to its high processing speed (14K sentences per second). The predicted likelihood of key tokens is averaged as the confidence score. Results in Fig. 9 verify that confidence and model performance are highly correlated.
Model setup. Since the CLIP pre-trained ViT-L/14 [clip] encodes each $14\times 14$ patch to a visual token, the resolution of the input image should be integer multiples of 14. Therefore, we first pad the size of input images to integer multiples of 14 with zero-padding. We encode the image patches into visual tokens using the CLIP pre-trained ViT-L/14 [clip], with each token having a channel of 1024. The vision abstractor can reduce the number of vision tokens to 64 and map the vision tokens to the hidden dimension of the LLM, which is 4096. Without the vision abstractor, the maximum resolution is limited to 672, constrained by computation resources (RTX A6000 GPUs). However, with the vision abstractor, we can process images with much larger resolutions (up to $2500\times 2500$). In our experiments, the maximum image resolution is $1092\times 1456$, thus the resolutions of all images are retained. The vision abstractor consists of four transformer layers with 64 learnable query embeddings. In LoRA of LLM, the parameters of rank and scale factor are both set as 16. The projection weights in each attention layer are fine-tuned using LoRA technique.
Training and inference setup. DepictQA-Wild is trained for 1 epoch with batch size 64. Adam optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, weight decay 0.001, and learning rate 0.0002 is used for training. During inference, the temperature is set to 0, since lots of predicted information (e.g., distortion, comparison result) need to be certain.
V Experiments
V-A Metrics and Baselines
@@
Accuracy, SRCC, and PLCC. The accuracy metric is utilized for distortion identification and instant rating tasks. VLMs usually produce diverse textual outputs, and we transform them into brief results for accuracy calculation. Specifically, we prompt our model with “Answer the question using a single word or phrase” to encourage direct output of brief responses. For baseline models, we include all potential answers in the prompt and instruct the model to identify the most accurate one. We emphasize that our key motivation is to generate descriptive language rather than quality scores. However, our approach can produce quality scores using pair-wise comparison if required. The quality scores are assessed using Spearman Rank Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC).
@@
GPT-4 score, BLEU, and ROUGE-L. We employ the GPT-4 score to evaluate assessment reasoning and comparison reasoning tasks, following [llava]. Specifically, we provide GPT-4 with both the model-generated response and the corresponding ground truth response. GPT-4 assesses the helpfulness, relevance, accuracy, and level of detail in the model-generated response relative to the ground truth, assigning an overall score on a scale of 0 to 10, where a higher score indicates better quality. This average score is subsequently normalized to a scale of 0 to 100%, reported as the GPT-4 score metric. We further evaluate the reasoning tasks with classical metrics including BLEU and ROUGE-L score following [metric_rouge, attention].
@@
Baselines. We categorize our baseline methods into general-purpose VLMs and IQA-specific VLMs. For general VLMs, we include mPLUG-Owl2 [mplug-owl2], LLaVA-1.6 [llavanext], LLaVA-OneVision-1.5 [llava_ov_1.5], InternVL2.5 [internvl_2.5], Qwen2.5-VL [qwen2.5_vl], and the proprietary GPT-4V [gpt4v]. We use 7B/8B version for open-source VLMs for fair comparison. IQA-specific VLMs are represented by Q-Instruct [qinstruct], Co-Instruct [coinstruct], CLIP-like LIQE [liqe], and reasoning-based IQA model Q-Insight [qinsight]. Note that Q-Instruct only supports single-image inputs, thus we only test it on non-reference single-image assessment tasks. Additionally, we compare traditional score-based IQA methods including full-reference ones (PSNR, SSIM [ssim], LPIPS [lpips], DISTS [dists]) and non-reference ones (NIQE [niqe], ClipIQA [clipiqa], MUSIQ [musiq], MANIQA [maniqa]) in instant rating task and score regression experiments.
V-B Results on Benchmarks
@@
TABLE III: Distortion identification results under both single-distortion and multi-distortion cases. The accuracy metric is reported in the full-reference / non-reference settings. The best results are highlighted in bold. Our DepictQA-Wild greatly outperforms all baselines and maintains its high accuracy in out-of-distribution (OOD) setting.
@@
mPLUG-Owl2 LLaVA-1.6 LLaVA-OV-1.5 InternVL2.5 Qwen2.5-VL GPT-4V Q-Instruct Co-Instruct LIQE Q-Insight Ours Ours-OOD
Single-dist. 10.1 / 11.6 14.0 / 15.3 23.9 / 30.4 18.7 / 20.3 33.0 / 31.8 46.7 / 45.2 -/ 15.5 27.2 / 34.4 -/ 33.1 48.4 / 49.9 97.7 / 94.1 82.1 / 73.2
Multi-dist. 10.8 / 10.7 12.0 / 12.1 32.6 / 38.3 35.0 / 36.8 36.5 / 38.3 41.5 / 39.8 -/ 23.9 30.2 / 33.3 -/ 31.4 43.5 / 49.6 91.3 / 89.3 76.6 / 77.2
@@
TABLE IV: Instant rating results on multiple benchmarks in the full-reference / non-reference setting with the accuracy metric. The best results are shown in bold. Q-Instruct is tested by inputting single images to calculate quality scores, and then compare the scores to rate. DepictQA-Wild surpasses all baselines by a large margin.
[…]
Full-reference Score-based IQA PSNR 68.9 / 78.7 / 80.9 / 85.0 / 89.7 / 78.0 / 80.2 /
SSIM 69.7 / 77.1 / 82.6 / 78.7 / 88.1 / 76.8 / 78.8 /
LPIPS 79.4 / 79.7 / 84.2 / 86.6 / 91.3 / 85.4 / 84.4 /
DISTS 79.7 / 85.8 / 84.6 / 87.0 / 93.1 / 88.5 / 86.5 /
Non-reference Score-based IQA NIQE / 49.9 / 66.9 / 59.7 / 65.0 / 86.9 / 82.2 / 68.4
ClipIQA / 59.7 / 75.8 / 72.6 / 85.8 / 65.8 / 47.0 / 67.8
MUSIQ / 59.2 / 76.1 / 77.8 / 80.1 / 87.2 / 81.1 / 76.9
MANIQA / 54.9 / 68.4 / 79.2 / 77.3 / 75.4 / 63.5 / 69.8
General VLM mPLUG-Owl2 50.1 / 50.1 50.6 / 50.8 49.6 / 49.6 48.6 / 48.5 49.9 / 50.1 50.6 / 50.5 49.9 / 49.9
LLaVA-1.6 54.1 / 56.2 50.4 / 51.9 52.0 / 52.6 54.2 / 57.0 54.4 / 56.5 54.3 / 53.1 53.2 / 54.6
LLaVA-OV-1.5 50.0 / 49.7 50.1 / 51.4 50.2 / 49.6 51.9 / 49.8 51.3 / 47.8 50.3 / 52.3 50.6 / 50.1
InternVL2.5 50.5 / 62.8 50.2 / 75.5 50.3 / 67.0 52.4 / 83.9 52.0 / 73.0 49.4 / 73.2 50.8 / 72.6
Qwen2.5-VL 49.6 / 55.3 55.8 / 76.2 56.7 / 67.9 61.9 / 83.3 54.5 / 64.3 50.8 / 55.6 54.9 / 67.1
GPT-4V 70.3 / 63.2 83.2 / 81.5 78.5 / 78.2 84.4 / 88.1 79.6 / 72.7 70.6 / 67.6 77.8 / 75.2
IQA-specific VLM Q-Instruct - / 41.6 - / 81.7 - / 74.6 - / 88.8 - / 73.1 - / 48.5 - / 68.1
Co-Instruct 49.8 / 50.7 52.0 / 82.4 50.6 / 72.5 59.3 / 85.0 50.0 / 70.3 50.0 / 58.0 52.0 / 69.8
Q-Insight 49.9 / 50.8 51.1 / 85.0 63.4 / 68.6 54.2 / 90.8 48.9 / 50.7 55.8 / 58.0 53.9 / 67.3
DepictQA-Wild 84.7 / 82.4 93.6 / 93.1 90.5 / 90.0 96.9 / 96.4 92.1 / 91.8 90.0 / 89.6 91.3 / 90.6
Quantitative results of distortion identification are shown in Sec. V-B. First, for open-source general VLMs and Co-Instruct, the performance in the non-reference setting is generally better than in the full-reference setting. This is counterintuitive, as humans typically find full-reference comparison easier. The reason is that current VLMs are primarily trained without reference, thus it is hard to use the reference image. Second, for open-source general VLMs, the performance in the multi-distortion case is higher than in the single-distortion case. In fact, for multi-distortion samples like “blur, darken”, when the model predicts only one distortion “blur”, it still achieves 0.5 accuracy. Therefore, for multi-distortion cases, achieving very high accuracy (>>90%) is difficult, whereas obtaining moderate accuracy (<<40%) is comparatively easier than in single-distortion cases. Third, the proprietary GPT-4V outperforms other general-purpose VLMs and exceeds specialized IQA VLMs like Q-Instruct and Co-Instruct, achieving similar results with reasoning-based IQA model Q-Insight. Fourth, DepictQA-Wild stably surpasses all baseline methods, demonstrating our model’s efficacy. Finally, we evaluate our model in an out-of-distribution (OOD) setting, i.e., for a particular category of distortion (e.g., noise), we use some sub-categories (e.g., Gaussian noise) during training, and other sub-categories (e.g., impulse noise) for evaluation. Results in the last column of Sec. V-B show that our method maintains high accuracy even under such an OOD setting.
Quantitative results of instant rating are demonstrated in Sec. V-B. First, in the full-reference context, traditional score-based methods, even the simplest PSNR, outperform all general VLMs including GPT-4V and prior IQA-specific VLMs, indicating the inadequacy of existing VLMs in full-reference IQA tasks. Second, conversely, in the non-reference scenario, GPT-4V and Co-Instruct excel beyond most score-based approaches, except MUSIQ. Third, for some general VLMs (i.e., InternVL2.5 and Qwen2.5-VL) and IQA-specific VLM Co-Instruct, the performance in the non-reference setting consistently surpasses that in the full-reference setting, as these models struggle to effectively utilize reference images even with explicit prompts. This further demonstrates the necessity of unifying full-reference and non-reference settings. Fourth, the reasoning-based Q-Insight model performs much better in some cases (e.g., non-reference KADID and TID2013) than in others (e.g., LIVE-MD and MDID2013), indicating its instability across datasets. Finally, DepictQA-Wild demonstrates superior performance across both settings by a large margin, showcasing its substantial advantage.